# Efficient semantic uncertainty quantification in language models via diversity-steered sampling

### v1

@@ 2 Estimating the predictive uncertainty of free-form responses
Let $x$ be an input (e.g. a question) and $\theta$ the weights of a pretrained language model. We wish to measure the total predictive uncertainty of the model’s output under semantic clustering. Following Aichberger et al. [4], we define the distribution over semantic clusters $c\in\mathcal{C}$ by marginalizing over all output sequences $y$:
@@ 3.1 Diversity-steered sampling
To compute similarity, we construct an intermediate input $z^{(t)}$ by substituting the current proposal $y^{(t)}$ into the masked positions of $y^{(t+1)}$, and then evaluate:
$\log q(y^{(t)}|y^{(t+1)})=\log p(y^{(t)}|y^{(t+1)})-\lambda\max_{s\in\mathcal{S}}E(z^{(t)},s).$ (7)
@@ 3.2 Fine-tuning NLI for partial sequences
Standard NLI models expect full premise–hypothesis pairs, but our sampler requires entailment scores for partially generated or masked text. To adapt an off‐the‐shelf NLI model with minimal overhead, we begin by loading a model finetuned on a natural language understanding dataset such as the MNLI benchmark dataset [11] (e.g. DeBERTa‐large‐MNLI [12]) and freezing all of its existing parameters. We then consider a special token: [TRUNC] [^2] If the token does not exist in the tokenizer’s vocabulary, we add it and initialize its embedding randomly. in the case of ARMs and [MASK] in the case of MDMs. The token embedding and the model’s final classification layer are the only components allowed to update during finetuning.
@@ 3.3 Importance-reweighted estimators of uncertainty
We use the MI proxy introduced by Yadkori et al. [2, Algorithm 2], restricted here to the pairwise case n=2n{=}2, as a worked example of how one can quantify epistemic uncertainty with our semantically–diverse sample set. Rather than a faithful reproduction of the original MI implementation, we aim at demonstrating our diversity sampling within an importance sampling framework. The MI proxy operates on the assumption that successive answers are conditionally independent under the ground-truth distribution $p^{*}$:
@@
We draw $N$ pairs $z^{(i)}\coloneqq(s^{(i)}_{1},s^{(i)}_{2})$ from the diversity-steered distribution $q^{e}$ by iterative prompting. Each response token sequence is mapped to a semantic cluster $c^{(i)}_{j}{\in}\mathcal{C}$ using the NLI-reported bidirectional entailment score [1], rather than using the F1 score [14] as originally done in Yadkori et al. [2]. We find cluster centers $\mathcal{S}$ defined by all tuples $j,k$ such that $E(z^{(j)},z^{(k)})<\tau$ and each cluster center $z^{\star}\in\mathcal{S}$ is associated with a cluster $c(z^{\star})\coloneqq\{z^{\prime}:E(z^{\star},z^{\prime})\geq\tau\}$.
@@ 4 Related Work
Common approaches for uncertainty quantification in vision and classification tasks include Monte Carlo dropout [15], deep ensembles [16], and prior networks [17, 18]. While these methods have also been adapted for text classification and regression in NLP [19, 20, 21, 22], their extension to free-form NLG presents specific challenges due to semantic invariances inherent in generated text [1, 23]. Several studies have proposed prompting or fine-tuning language models to explicitly articulate their confidence levels [24, 25, 26, 27, 28, 29, 30, 31], though this typically requires additional supervision. Attention values may contain information about relevance or confidence [32, 33]. Alternatively, cross-examination leverages a secondary language model to evaluate uncertainty in another model’s outputs [34]. Predictive entropy quantifies the token-level entropy of the predictive distribution [17]. Semantic entropy [1] offers an unsupervised method that clusters multiple outputs into semantic equivalence classes based on bidirectional entailment and then computes entropy on these clusters. Complementary strategies, such as conformal prediction, can provide bounds on hallucinations under stronger theoretical assumptions [35, 36].
@@
Sampling heuristics aimed at enhancing output diversity, such as temperature scaling, top-$k$, and nucleus sampling [3] do not account for semantics. Diverse beam search (DBS) introduces diversity heuristics within beam search optimization [39]. Contrastive decoding enlists a secondary, weaker language model whose output tokens are penalized to encourage diverse token selection by the primary model [40]. Cluster-based beam search methods apply semantic clustering to prune beam candidates and diversify subsequent selections [41, 42], but this heavily depends on initial candidate diversity. Semantically Diverse Language Generation (SDLG) substitutes the most informative token in a fully generated sample and allows standard sampling to proceed from that token onward [4]. It does not, however, explicitly account for diversity within a running sample set and, because the token scoring involves a gradient of the NLI loss with respect to the NLI token space, requires additional implementation [e.g., 43] if the base LLM does not share the tokenization scheme or vocabulary with the NLI. Our approach differs by integrating a continuous semantic penalty directly into the sampling logits during text generation in a gradient-free manner. For comparative analyses of diversity-enhancing decoding strategies, see Ippolito et al. [44].
@@
Grouping model outputs based on semantic equivalence often reduces to paraphrase detection. Casting semantic equivalence as bidirectional entailment dates back to early linguistics work [45] and was later adopted in NLP [46, 47]. Early methods relied on lexical overlap [48, 49] or vector embedding similarities [50, 51]. Recent systems employ BERT-style encoders to build binary “paraphrases” vs. “not paraphrases” classifiers [12, 52, 53]. The entailment probability from NLI models has been used to cluster LLM generations into meaning-equivalent sets, enabling unsupervised uncertainty estimation [1, 4, 54].
@@ 5 Experiments
We evaluate our diversity‑promoting sampling scheme for estimating SE [1] and MI [2]. The target estimators are the importance-reweighted versions with control variates: $\hat{H}_{\rm cv}$ in Equation 14 and $\hat{I}_{\rm cv}$ in Equation 15, respectively. Due to space constraints, implementation details, additional results for $\hat{H}$ including ablation studies, and results for $\hat{I}_{\rm cv}$ are deferred to Appendices˜B, C and D. As observed by Aichberger et al. [4], it is common to employ diversity heuristics during sampling without correcting for the introduced bias with principled importance reweighting [e.g., 1, 2]. Although one of our contributions is the variance-reduced application of importance correction with control variates, we apply the same estimation procedure to the final generations from all sampling schemes, to enable a fair comparison. Estimation is performed on top of the semantic clusters created using Kuhn et al. [1, Algorithm 1]. In brief, clustering works by querying the DeBERTa-large-MNLI model [12], finetuned on the NLI dataset MNLI [11], on every pair of sampled responses. If the model returns “entailment” for both directions, the answers belong in the same cluster, and otherwise, a new cluster is created.
@@ Appendix A Extension to masked diffusion models
9:     for each candidate token $\mathrm{next}\in\mathcal{V}$ do $\triangleright$ Can alternatively consider the top-$k$ tokens only
@@ B.3 Adaptively tuning the diversity parameter $\lambda$
with sequence-level learning rate $\eta_{\text{seq}}$. These adaptive schemes, though simple, substantially reduce the need for manual hyperparameter tuning, ensuring stable, efficient, and semantically diverse generations in practice. While $\lambda$ can similarly be calibrated on a held-out dataset using $E_{\rm target}$ or $V_{\rm target}$ as target values [e.g., 2], we do not opt for a separate calibration step, as this expends additional compute and it is difficult to curate an IID calibration dataset in practice.

### v2

@@ 2 Estimating the predictive uncertainty of free-form responses
Let $x$ be an input (e.g., a question) and $\theta$ the weights of a pretrained language model. We wish to measure the total predictive uncertainty of the model’s output under semantic clustering. Following Aichberger et al. [3], we define the distribution over semantic clusters $c\in\mathcal{C}$ by marginalizing over all output sequences $y$:
@@ 3.1 Diversity-steered sampling
To compute similarity, we construct an intermediate input $z^{(t)}$ by substituting the current proposal $y^{(t)}$ into the masked positions of $y^{(t+1)}$, and then evaluate the unnormalized distribution:
$\log\tilde{q}(y^{(t)}\ |\ y^{(t+1)})=\log p(y^{(t)}\ |\ y^{(t+1)})-\lambda\max_{s\in\mathcal{S}}E(z^{(t)},s).$ (7)
@@ 3.2 Fine-tuning NLI for partial sequences
Standard NLI models expect full premise–hypothesis pairs, but our sampler requires entailment scores for partially generated or masked text. To adapt an off‐the‐shelf NLI model with minimal overhead, we begin by loading a model finetuned on a natural language understanding dataset such as the MNLI benchmark dataset [62] (e.g., DeBERTa‐large‐MNLI [20]) and freezing all of its existing parameters. We then consider a special token: [TRUNC] [^2] If the token does not exist in the tokenizer’s vocabulary, we add it and initialize its embedding randomly. in the case of ARMs and [MASK] in the case of MDMs. The token embedding and the model’s final classification layer are the only components allowed to update during finetuning.
@@ 3.3 Importance-reweighted estimators of uncertainty
We use the MI proxy introduced by Yadkori et al. [63, Algorithm 2], restricted here to the pairwise case n=2n{=}2, as a worked example of how one can quantify epistemic uncertainty with our semantically–diverse sample set. Rather than a faithful reproduction of the original MI implementation, we aim at demonstrating our diversity sampling within an importance sampling framework. The MI proxy operates on the assumption that successive answers are conditionally independent under the ground-truth distribution $p^{*}$:
@@
We draw $N$ pairs $z^{(i)}\coloneqq(s^{(i)}_{1},s^{(i)}_{2})$ from the diversity-steered distribution $q^{e}$ by iterative prompting. Each response token sequence is mapped to a semantic cluster $c^{(i)}_{j}{\in}\mathcal{C}$ using the NLI-reported bidirectional entailment score [28], rather than using the F1 score as originally done in Yadkori et al. [63]. We find cluster centers $\mathcal{S}$ defined by all tuples $j,k$ such that $E(z^{(j)},z^{(k)})<\tau$ and each cluster center $z^{\star}\in\mathcal{S}$ is associated with a cluster $c(z^{\star})\coloneqq\{z^{\prime}:E(z^{\star},z^{\prime})\geq\tau\}$.
@@ 4 Related Work
Common approaches for uncertainty quantification in vision and classification tasks include Monte Carlo dropout [15], deep ensembles [30], and prior networks [38, 39]. While these methods have also been adapted for text classification and regression in NLP [23, 12, 60, 61], their extension to free-form NLG presents specific challenges due to semantic invariances inherent in generated text [17, 28]. Several studies have proposed prompting or fine-tuning language models to explicitly articulate their confidence levels [26, 40, 9, 16, 49, 55, 19], though this typically requires additional supervision. Attention values may contain information about relevance or confidence [35, 14]. Alternatively, cross-examination leverages a secondary language model to evaluate uncertainty in another model’s outputs [10]. Predictive entropy quantifies the token-level entropy of the predictive distribution [38]. Semantic entropy [28] offers an unsupervised method that clusters multiple outputs into semantic equivalence classes based on bidirectional entailment and then computes entropy on these clusters. Complementary strategies using the conformal framework can provide bounds on errors under stronger theoretical assumptions [47, 5].
@@
Sampling heuristics aimed at enhancing output diversity, such as temperature scaling, top-$k$, and nucleus sampling [21] do not account for semantics. Diverse beam search (DBS) introduces diversity heuristics within beam search optimization [57]. Contrastive decoding enlists a secondary, weaker language model whose output tokens are penalized to encourage diverse token selection by the primary model [32]. Cluster-based beam search methods apply semantic clustering to prune beam candidates and diversify subsequent selections [31, 54], but this heavily depends on initial candidate diversity. Semantically Diverse Language Generation (SDLG) substitutes the most informative token in a fully generated sample and allows standard sampling to proceed from that token onward [3]. It does not, however, explicitly account for diversity within a running sample set and, because the token scoring involves a gradient of the NLI loss with respect to the NLI token space, requires additional implementation [e.g., 6] if the base LLM does not share the tokenization scheme or vocabulary with the NLI. Our approach differs by integrating a continuous semantic penalty directly into the sampling logits during text generation in a gradient-free manner. For comparative analyses of diversity-enhancing decoding strategies, see Ippolito et al. [22].
@@
Grouping model outputs based on semantic equivalence often reduces to paraphrase detection. Casting semantic equivalence as bidirectional entailment dates back to early linguistics work [11] and was later adopted in NLP [44, 4, 45]. Early methods relied on lexical overlap [46] or vector embedding similarities [53, 64]. [8] explored the use of bilingual parallel corpora for paraphrase extraction and ranking. BERT-style encoders can be used to build binary “paraphrases” vs. “not paraphrases” classifiers [13, 59]. The entailment probability from NLI models has been used to cluster LLM generations into meaning-equivalent sets, enabling unsupervised uncertainty estimation [28, 3, 34].
@@ 5 Experiments
We evaluate our diversity‑promoting sampling scheme for estimating SE [28] and MI [63]. The target estimators are the importance-reweighted versions with control variates: $\hat{H}_{\rm cv}$ in Equation 14 and $\hat{I}_{\rm cv}$ in Equation 15, respectively. Due to space constraints, implementation details, additional results for $\hat{H}$ including ablation studies, and results for $\hat{I}_{\rm cv}$ are deferred to Appendices˜B, C and D. As observed by Aichberger et al. [3], it is common to employ diversity heuristics during sampling without correcting for the introduced bias with principled importance reweighting [e.g., 28, 63]. Although one of our contributions is the variance-reduced application of importance correction with control variates, we apply the same estimation procedure to the final generations from all sampling schemes, to enable a fair comparison. Estimation is performed on top of the semantic clusters created using Kuhn et al. [28, Algorithm 1]. In brief, clustering works by querying the DeBERTa-large-MNLI model [20], finetuned on the NLI dataset MNLI [62], on every pair of sampled responses. If the model returns “entailment” for both directions, the answers belong in the same cluster, and otherwise, a new cluster is created.
@@ Appendix A Extension to masked diffusion models
9:     for each candidate token $\mathrm{next}\in\mathcal{V}$ do $\triangleright$ Alternatively consider the top-$k$ tokens only
@@ B.3 Adaptively tuning the diversity parameter $\lambda$
with sequence-level learning rate $\eta_{\text{seq}}$. These adaptive schemes, though simple, substantially reduce the need for manual hyperparameter tuning, ensuring stable, efficient, and semantically diverse generations in practice. While $\lambda$ can similarly be calibrated on a held-out dataset using $E_{\rm target}$ or $V_{\rm target}$ as target values [e.g., 63], we do not opt for a separate calibration step, as this expends additional compute and it is difficult to curate an IID calibration dataset in practice.